a fair coin is tossed four times, let x be the product of the number of heads times the number of tails...
Find the most likely value of X

Answer by stanbon(75887)

a fair coin is tossed four times, let x be the product of the number of heads times the number of tails...
Find the most likely value of X
-----
Binomial Problem with n = 4 and P(heads) = 1/2
----
Expected value = mean = np = (1/2)4 = 2
-----
Cheers,
Stan H.
============